A 0.12-kg ball is moving at 6 m/s when it is hit by a bat, causing it to reverse direction and have a speed of 14 m/s. What is the change in the magnitude of the momentum of the ball?
1
Expert's answer
2020-06-08T10:30:06-0400

The initial momentum of the ball, by definition, is the following:


"p_1 = 0.12 kg\cdot 6m/s = 0.72 kg\cdot m/s"

After the collision, the momentum will be (sign minus due to the direction reversing):


"p_2 = -0.12kg\cdot 14 m/s = -1.68 kg\cdot m/s"

Thus, the change in the magnitude of the momentum will be:


"\Delta p = |p_2 - p_1| = |-1.68 - 0.72 | = 2.4 kg\cdot m/s"

Answer. The change in the magnitude of the momentum is 2.4 kg*m/s.